OpenStudy (anonymous):

∫dx/x2+x+1 and ∫dx/(x^2 -x) are my questions. Thanks for helping.

OpenStudy (anonymous):

lajeik? is that u

OpenStudy (anonymous):

lajeik fleming

OpenStudy (anonymous):

Lajeik Fleming? the little black sleek-feathered one

OpenStudy (anonymous):

omfg its lajeik

OpenStudy (anonymous):

Have you tried partial fractions on this one?

OpenStudy (anonymous):

\[(x^2 - x + 2) / ((x-1)(x^2 + x + 1) ) = A/(x+1) + (Bx + C)/(x^2 + x +1) \] No I haven't. Is this how I begin?

OpenStudy (anonymous):

Yes, your decomposition so far is good. All that's left is finding A, B, and C.

OpenStudy (anonymous):

ok thank you I have found A B and C which leads me to integrate \[(x-4)/(x^2 + x + 1)\] How do I integrate this?

OpenStudy (anonymous):

So, you're integrating this \[\frac{x-4}{x^2+x+1}\] or \[\frac{x^2-x+2}{(x-1)(x^2+x+1)}?\] Sorry, I'm a bit confused.

OpenStudy (anonymous):

after partial fractions I am left with a longer integral of \[2/3 \int\limits_{} 1/(x+1) dx +1/3 \int\limits_{} (x-4)/(x^2 + x + 1) dx\] How do I start the second integral ?

OpenStudy (anonymous):

should be 1/(x-1) in the first part..

OpenStudy (anonymous):

Ah, thanks. \[\int\frac{x-4}{x^2+x+1}dx\] Notice that the derivative of the denominator is 2x + 1. Rewriting the numerator, you get \[\frac{1}{2}\int\frac{2x-8}{x^2+x+1}dx\\ \frac{1}{2}\int\frac{2x+1-9}{x^2+x+1}dx\\ \frac{1}{2}\left[\int\frac{2x+1}{x^2+x+1}dx-\frac{1}{2}\int\frac{9}{x^2+x+1}dx\right] \]

OpenStudy (anonymous):

Sorry, the last line should be \[\frac{1}{2}\left[\int\frac{2x+1}{x^2+x+1}dx-\int\frac{9}{x^2+x+1}dx\right]\]

OpenStudy (anonymous):

Now, making the substitution for the first integral: \[u=x^2+x+1\\ du=(2x+1)\;dx\] \[\frac{1}{2}\left[\int\frac{1}{u}du-\int\frac{9}{x^2+x+1}dx\right]\] For the second integral, complete the square in the denominator.

OpenStudy (anonymous):

Okay. We haven't learned this in class and I don't think it will be on the test. I thank you so much for explaining this much so far. What math is this taught in? I'm in Calc 272 at a community college and I'm wondering when I have to know this. lol.. Thanks so much for your help.. wow.. unbelievable I haven't been taught this yet.

OpenStudy (anonymous):

There isn't really an explicit name for this method. Just some algebraic manipulation to get what you want.

OpenStudy (anonymous):

What math did you learn this in?

OpenStudy (anonymous):

BC calc, two years ago

OpenStudy (anonymous):

Thank you so much for your help!

OpenStudy (anonymous):

You're welcome. If you're curious, the last integral (with the completing the square) later uses trig substitution. If you've learned that, it might be good practice.

OpenStudy (anonymous):

Sure I'll try it.

OpenStudy (anonymous):

∫dx/x2+x+1

OpenStudy (anonymous):

∫dx/(x^2 -x)

OpenStudy (anonymous):

∫dx/x2+x+1 and ∫dx/(x^2 -x) are my questions. Thanks for helping.